Tara earned the following quiz scores in her biology class:92,76,80,83 and 91. What must she get on her next quiz to have an average grade of 85% in the class?
A.88
B.3
C.82

Answer by dabanfield(803)

Tara earned the following quiz scores in her biology class:92,76,80,83 and 91. What must she get on her next quiz to have an average grade of 85% in the class?
Let x be the score on the next quiz:
Then the average on the 6 tests will be:
(92+76+80+83+91+x)/6
Set this equal to 85 and solve for x:
(92+76+80+83+91+x)/6 = 85
Multiply both sides by 6:
92+76+80+83+91+x= 6*85 = 510
x = 510-92-76-80-83-91
Finish the calculation.